\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{10}

\bibitem{BeelenDoumen}
P.~H.~T. Beelen and J.~M. Doumen, \emph{Pseudorandom sequences from elliptic
  curves}, Finite fields with applications to coding theory, cryptography and
  related areas ({O}axaca, 2001), Springer, Berlin, 2002, pp.~37--52.

\bibitem{cantor}
David~G. Cantor, \emph{Computing in the {J}acobian of a hyperelliptic curve},
  Math. Comp. \textbf{48} (1987), no.~177, 95--101.

\bibitem{ChenECLegendre}
Zhixiong Chen, \emph{Elliptic curve analogue of {L}egendre sequences}, Monatsh.
  Math. \textbf{154} (2008), no.~1, 1--10.

\bibitem{cohen2005handbook}
Henri Cohen, Gerhard Frey, Roberto Avanzi, Christophe Doche, Tanja Lange, Kim
  Nguyen, and Frederik Vercauteren (eds.), \emph{Handbook of elliptic and
  hyperelliptic curve cryptography}, Discrete Mathematics and its Applications
  (Boca Raton), Chapman \& Hall/CRC, Boca Raton, FL, 2006.

\bibitem{ElMaShp2001}
Edwin El~Mahassni and Igor Shparlinski, \emph{On the uniformity of distribution
  of congruential generators over elliptic curves}, Sequences and their
  applications ({B}ergen, 2001), Discrete Math. Theor. Comput. Sci. (Lond.),
  Springer, London, 2002, pp.~257--264.

\bibitem{Rezaeian}
Reza~Rezaeian Farashahi, \emph{Extractors for {J}acobian of hyperelliptic
  curves of genus 2 in odd characteristic}, Cryptography and coding, Lecture
  Notes in Comput. Sci., vol. 4887, Springer, Berlin, 2007, pp.~313--335.

\bibitem{Reza2}
Reza~Rezaeian Farashahi and Ruud Pellikaan, \emph{The quadratic extension
  extractor for (hyper)elliptic curves in odd characteristic}, Arithmetic of
  finite fields, Lecture Notes in Comput. Sci., vol. 4547, Springer, Berlin,
  2007, pp.~219--236.

\bibitem{Galbraith-book}
Steven~D. Galbraith, \emph{Mathematics of public key cryptography}, Cambridge
  University Press, Cambridge, 2012.

\bibitem{GongBersonStinson99}
Guang Gong, Thomas~A. Berson, and Douglas~R. Stinson, \emph{Elliptic curve
  pseudorandom sequence generators}, Selected areas in cryptography
  ({K}ingston, {ON}, 1999), Lecture Notes in Comput. Sci., vol. 1758, Springer,
  Berlin, 2000, pp.~34--48.

\bibitem{GongLam2001}
Guang Gong and Charles C.~Y. Lam, \emph{Linear recursive sequences over
  elliptic curves}, Sequences and their applications ({B}ergen, 2001), Discrete
  Math. Theor. Comput. Sci. (Lond.), Springer, London, 2002, pp.~182--196.

\bibitem{grant1990formal}
David Grant, \emph{Formal groups in genus two}, J. Reine Angew. Math.
  \textbf{411} (1990), 96--121.

\bibitem{hallgren}
Sean Hallgren, \emph{Linear congruential generators over elliptic curves},
  Tech. report, Dept. of Comp. Sci., Cornegie Mellon Univ., 1994.

\bibitem{hartshorneAlgebraicGeometry1977}
Robin Hartshorne, \emph{Algebraic geometry}, Springer-Verlag, New
  York-Heidelberg, 1977, Graduate Texts in Mathematics, No. 52.

\bibitem{hessLinearComplexityMultidimensional2005}
Florian Hess and Igor~E. Shparlinski, \emph{On the linear complexity and
  multidimensional distribution of congruential generators over elliptic
  curves}, Des. Codes Cryptogr. \textbf{35} (2005), no.~1, 111--117.

\bibitem{hindryDiophantineGeometryIntroduction2000}
Marc Hindry and Joseph~H. Silverman, \emph{Diophantine geometry}, Graduate
  Texts in Mathematics, vol. 201, Springer-Verlag, New York, 2000, An
  introduction.

\bibitem{Koblitz-book}
Neal Koblitz, \emph{Algebraic aspects of cryptography}, Algorithms and
  Computation in Mathematics, vol.~3, Springer-Verlag, Berlin, 1998, With an
  appendix by Alfred J. Menezes, Yi-Hong Wu and Robert J. Zuccherato.
  \MR{1610535}

\bibitem{KohelShparlinski}
David~R. Kohel and Igor~E. Shparlinski, \emph{On exponential sums and group
  generators for elliptic curves over finite fields}, Algorithmic number theory
  ({L}eiden, 2000), Lecture Notes in Comput. Sci., vol. 1838, Springer, Berlin,
  2000, pp.~395--404.

\bibitem{langeFormulaeArithmeticGenus2005}
Tanja Lange, \emph{Formulae for arithmetic on genus 2 hyperelliptic curves},
  Appl. Algebra Engrg. Comm. Comput. \textbf{15} (2005), no.~5, 295--328.

\bibitem{ShparlinskiLange05}
Tanja Lange and Igor Shparlinski, \emph{Collisions in fast generation of ideal
  classes and points on hyperelliptic and elliptic curves}, Appl. Algebra
  Engrg. Comm. Comput. \textbf{15} (2005), no.~5, 329--337.

\bibitem{langeCertainExponentialSums2005}
Tanja Lange and Igor~E. Shparlinski, \emph{Certain exponential sums and random
  walks on elliptic curves}, Canad. J. Math. \textbf{57} (2005), no.~2,
  338--350.

\bibitem{MeidlWinterhof}
Wilfried Meidl and Arne Winterhof, \emph{Linear complexity of sequences and
  multisequences}, Handbook of Finite Fields, CRC Press, Boca Raton, 2013.

\bibitem{MeraiEcMult}
L\'{a}szl\'{o} M\'{e}rai, \emph{Construction of pseudorandom binary sequences
  over elliptic curves using multiplicative characters}, Publ. Math. Debrecen
  \textbf{80} (2012), no.~1-2, 199--213.

\bibitem{MeraiECAdditive}
\bysame, \emph{Remarks on pseudorandom binary sequences over elliptic curves},
  Fund. Inform. \textbf{114} (2012), no.~3-4, 301--308.

\bibitem{meraiLinearComplexityProfile2015}
L\'{a}szl\'{o} M\'{e}rai and Arne Winterhof, \emph{On the linear complexity
  profile of some sequences derived from elliptic curves}, Des. Codes Cryptogr.
  \textbf{81} (2016), no.~2, 259--267.

\bibitem{mumfordTataLecturesTheta2007a}
David Mumford, \emph{Tata lectures on theta. {II}}, Modern Birkh\"{a}user
  Classics, Birkh\"{a}user Boston, Inc., Boston, MA, 2007, Jacobian theta
  functions and differential equations, With the collaboration of C. Musili, M.
  Nori, E. Previato, M. Stillman and H. Umemura, Reprint of the 1984 original.

\bibitem{Nied1}
Harald Niederreiter, \emph{Linear complexity and related complexity measures
  for sequences}, Progress in cryptology---{INDOCRYPT} 2003, Lecture Notes in
  Comput. Sci., vol. 2904, Springer, Berlin, 2003, pp.~1--17.

\bibitem{schmid1995affine}
Joachim Schmid, \emph{On the affine {B}ezout inequality}, Manuscripta Math.
  \textbf{88} (1995), no.~2, 225--232.

\bibitem{survey_Shparlinski}
Igor~E. Shparlinski, \emph{Pseudorandom number generators from elliptic
  curves}, Recent trends in cryptography, Contemp. Math., vol. 477, Amer. Math.
  Soc., Providence, RI, 2009, pp.~121--141.

\bibitem{stichtenothAlgebraicFunctionFields2009}
Henning Stichtenoth, \emph{Algebraic function fields and codes}, second ed.,
  Graduate Texts in Mathematics, vol. 254, Springer-Verlag, Berlin, 2009.

\bibitem{TopWinterhof2007}
Alev Topuzo\u{g}lu and Arne Winterhof, \emph{Pseudorandom sequences}, Topics in
  geometry, coding theory and cryptography, Algebr. Appl., vol.~6, Springer,
  Dordrecht, 2007, pp.~135--166.

\bibitem{Winterhof2010}
Arne Winterhof, \emph{Linear complexity and related complexity measures},
  Selected topics in information and coding theory, Ser. Coding Theory
  Cryptol., vol.~7, World Sci. Publ., Hackensack, NJ, 2010, pp.~3--40.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath,amssymb,amsbsy,amsfonts,amsthm,latexsym,mathabx,
            amsopn,amstext,amsxtra,euscript,amscd,stmaryrd,mathrsfs,
            cite,array,mathtools,enumerate}

\usepackage{url}
\usepackage[colorlinks,linkcolor=blue,anchorcolor=blue,citecolor=blue,backref=page]{hyperref}

\usepackage{color}
\usepackage{a4wide}

\usepackage{color}
\usepackage{float}

\hypersetup{breaklinks=true}

\usepackage[english]{babel}

\usepackage{mathtools}
\usepackage{todonotes}

\usepackage[norefs,nocites]{refcheck}

\usepackage{enumitem}

\allowdisplaybreaks

\usepackage{todonotes}
\usepackage{url}
\usepackage[colorlinks,linkcolor=blue,anchorcolor=blue,citecolor=blue,backref=page]{hyperref}
\usepackage[noabbrev,capitalize]{cleveref}

\usepackage{tikz}
\usetikzlibrary{arrows.meta}
\usepackage{tikz-cd}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{prop}[theorem]{Proposition}

\numberwithin{equation}{section}
\numberwithin{theorem}{section}

 \newcommand{\F}{\mathbb{F}}
\newcommand{\K}{\mathbb{K}}

\newcommand{\U}{U}

\newcommand{\X}{Z}

\DeclareMathOperator{\Char}{char}

\newcommand{\Fq}{\mathbb{F}_q}
\newcommand{\Fqb}{\overline{\mathbb{F}}_q}

\def\cN{{\mathcal N}}
\def\cO{{\mathcal O}}

\title[Linear complexity of  hyperelliptic curve sequences]{Linear complexity of some sequences derived from hyperelliptic curves of genus 2}

\author[V.~Anupindi]{Vishnupriya Anupindi}
 \author[L.~M{\'e}rai]{L{\'a}szl{\'o} M{\'e}rai}
 \address{Johann Radon Institute for Computational and Applied Mathematics, Austrian Academy of Sciences and Institute of Financial Mathematics and Applied Number Theory,
 Johannes Kepler University,  Altenberger Stra\ss e 69, A-4040 Linz, Austria} 
 \email{vishnupriya.anupindi@oeaw.ac.at}
 \email{laszlo.merai@oeaw.ac.at}

\begin{document}

\begin{abstract}
For a given hyperelliptic curve $C$ over a finite field with Jacobian $J_C$, we consider the hyperelliptic analogue of the congruential generator defined by  $W_n=W_{n-1}+D$ for $n\geq 1$ and $D,W_0\in J_C$. We show that curves of genus 2 produce sequences with large linear complexity.
\end{abstract}

\keywords{elliptic curve, hyperelliptic curve, linear complexity, congruential generator}

\subjclass[2020]{11G05, 11G20, 11K45, 11T71 }
\maketitle

\section{Introduction}

Let $\Fq$ be a finite field with characteristic $p\geq 3$ and consider the hyperelliptic curve $C$ of genus $g\geq 1$ defined by
\begin{equation} \label{def:hec}
    C: Y^2=f(X), 
\end{equation}
where $f(X)\in \Fq[X]$ is a monic polynomial of degree $2g+1$.

For details on hyperelliptic curves, see \cite{cohen2005handbook,Galbraith-book, Koblitz-book}.

We denote the $\Fq$-rational points of $C$ by $C(\Fq)$, which are the solutions over $\Fq$ of the defining equation \eqref{def:hec} together with a point $\cO$ at infinity. 
By the Hasse-Weil bound \cite[Theorem~5.2.3]{stichtenothAlgebraicFunctionFields2009}, we have
\begin{equation}\label{eq:number_of_points}
    ||C(\Fq)| -(q+1) |\leq 2g q^{1/2}.
\end{equation}

Unlike elliptic curves (curves with genus $g=1$), the points of a hyperelliptic curve with higher genus ($g\geq 2$) do not form an additive group. However, one can define a group operation by introducing the \emph{Jacobian} $J_C$ of the curve $C$.    

For an affine point $P=(x,y)\in C$, we write $-P=(x,-y)$ (and $-\cO=\cO$). A \emph{divisor} $D$ of $C(\Fqb)$ is an element of the free abelian group over the points of $C(\Fqb)$, e.\ g.\ $D =\sum_{P\in C(\Fqb)} n_P P$  with $n_P\in \mathbb{Z}$ and $n_P = 0$ for almost all points $P$.

Then any element $D\in J_C$ of the Jacobian  can be uniquely represented as a \emph{reduced divisor}
$$
\psi(D)=P_1+\dots +P_r-r\cO,
$$
where $1\leq r \leq g$, $\cO$ is the point of $C$ at infinity, $P_1, \dots, P_r\in C$, $P_i\neq \cO, 1\leq i \leq r$ and $P_i\neq -P_j$, $1\leq i<j\leq r$. The element $D$ is said to be defined over $\Fq$ if the \emph{Frobenius endomorphism}, defined by $\sigma((x,y))=(x^q,y^q)$ permutes the set $\{P_1,\dots, P_r\}$. We use $J_C(\Fq)$ to denote the set of elements of $J_C$ which are defined over $\Fq$. It follows from the Hasse-Weil Theorem   \cite[Theorem 5.1.15 and 5.2.1]{stichtenothAlgebraicFunctionFields2009}, that
\begin{equation}\label{eq:size_of_jac}
    (q^{1/2}-1)^{2g}\leq |J_{C}(\Fq)|\leq (q^{1/2}+1)^{2g}.
\end{equation}
It is common to represent the elements of the Jacobian by the \emph{Mumford representation} \cite{mumfordTataLecturesTheta2007a}. Let $D \in J_C(\Fq)$, then the Mumford representation $\eta(D)$ is a pair $[u,v]$ of polynomials such that 
	\renewcommand{\theenumi}{\alph{enumi}}
	\begin{enumerate}
		\item \label{mum-a} $u$ is monic,
		\item $u$ divides $f-v^2$,
		\item $\deg(v)< \deg(u)\leq g$.
	\end{enumerate}
Let $\psi(D)=\sum_{i=1}^r P_i-r \cO$, where $P_1, \dots, P_r\in C$, $P_i\neq \cO, 1\leq i \leq r$ and $P_i\neq -P_j$, $1\leq i<j\leq r$. Put $P_i=(x_i,y_i)$. Then $u=\prod_{i=1}^r(X-x_i)$ and $v(x_i)=y_i$ is of the same multiplicity as $P_i$ in $\psi(D)$.

One can define a group operation, denoted by $+$, on the Jacobian $J_C$. 
In Mumford coordinates, the group operation can be computed using Cantor's algorithm~\cite{cantor}. This algorithm can be made highly effective for small genus, which is the most important case in cryptographic applications.  See for example \cite{langeFormulaeArithmeticGenus2005} for the case of genus $g=2$.

\bigskip

In this paper, we study the properties of pseudorandomness of certain walks on the Jacobian $J_C(\Fq)$. Namely, let $D\in J_C(\Fq)$ and define the sequence
\begin{equation}\label{eq:S}
    W_n=D + W_{n-1}=nD + W_0, \quad n=1,2,\dots,
\end{equation}
with some initial value $W_0\in J_C(\Fq)$.

In the special case $g=1$, that is, when $C$ is an elliptic curve and $J_C\cong C$, the sequence $(W_n)$ has been suggested as a pseudorandom number generator in \cite{hallgren} and later many pseudorandom properties of this sequence have been studied \cite{BeelenDoumen,ElMaShp2001,GongBersonStinson99,GongLam2001,meraiLinearComplexityProfile2015,hessLinearComplexityMultidimensional2005,TopWinterhof2007,MeraiECAdditive,MeraiEcMult,ChenECLegendre}. 

In particular, the linear complexity of the coordinates of $(W_n)$ has been studied \cite{hessLinearComplexityMultidimensional2005, meraiLinearComplexityProfile2015,TopWinterhof2007}.
%\commL{$w\rightarrow a$ here}
We recall, that the \textit{$N$-th linear complexity} $L(s_n,N)$ of a sequence $(s_n)$ over the finite field $\Fq$ is defined as the smallest non-negative integer $ L $ such that the first $ N $ terms of the sequence $ (s_n) $ can be generated by a linear recurrence relation over $ \Fq $ of order $ L $, i.e.\ there exist $ c_0, c_1, \dots , c_{L-1} \in \Fq $ such that 
\begin{equation*}
	s_{n+L} = c_0s_{n} + c_1s_{n+1} + \dots  + c_{L-1}s_{n+L-1},\quad  0 \leq n \leq N-L-1.
\end{equation*}
The $N$-th linear complexity measures the unpredictability of a sequence and thus is an important figure of merit in cryptography. Clearly, large linear complexity is a desired (but not sufficient) property for such applications.  For more details, see \cite{MeidlWinterhof,Nied1, Winterhof2010}.

Hess and Shparlinski \cite{hessLinearComplexityMultidimensional2005} estimated the linear complexity of the coordinates for the elliptic curve analog of \eqref{eq:S}. Namely, let $x(\cdot)$ and $y(\cdot)$ be the coordinate functions of the curve such that for any affine point $P=(x(P),y(P))$. Among others, Hess and Shparlinski \cite{hessLinearComplexityMultidimensional2005} proved
$$
L(x(W_n ),N)\geq \min\left\{\frac{N}{4},\frac{t}{3}\right\},
$$
where $t$ is order of $D$, see also \cite{meraiLinearComplexityProfile2015,TopWinterhof2007}. 

\bigskip

In this paper, we estimate the linear complexity of the Mumford coordinates of the sequence $(W_n)$ defined by \eqref{eq:S} for hyperelliptic curves of genus $g=2$. More precisely, for $D\in J_C(\Fq)$, let $\eta(D)=[u(D),v(D)]$ be its Mumford representation and write
\begin{equation}\label{eq:coordinates}
    u(D)=u_2(D)X^2+u_1(D)X+u_0(D) \quad \text{and} \quad      v(D)=v_1(D)X+v_0(D), 
\end{equation}
where $u_2(D),u_1(D),u_0(D),v_1(D),v_0(D) \in\Fq$. By    \eqref{mum-a}, $u(D)$ is monic and thus $u_2(D)=1$  for all but approximately $q$ values $D$ and otherwise $u_2(D) = 0$. Thus, one cannot expect strong randomness properties of it. 
However, for the other coefficients, our result implies the lower bound for the linear complexity
\begin{equation}\label{eq:example}
L(u_0(W_n),N)\geq \left\lfloor c
		    \frac{\min\{t,N\}}{  q }\right\rfloor, 
\end{equation}
for some absolute and explicit constant $c>0$, where $t$ is the order of $D$, see Theorem~\ref{thm:main}. A similar bound holds for the other coefficients in \eqref{eq:coordinates} (except for $u_2$).

The most promising case is when the Jacobian $J_C(\Fq)$ is close to being a cyclic group, and $D$ has order $t=q^{2+o(1)}$ (cf. \eqref{eq:size_of_jac}).

\bigskip

In Section~\ref{sec:HEC}, we recall and prove the necessary tools concerning the arithmetic of hyperelliptic curves. In particular, we recall the Grant representation \cite{grant1990formal} of the Jacobian of hyperelliptic curves of genus $g=2$ as it gives explicit formulas for the addition law. In Section~\ref{sec:preliminaries}, we recall other required preliminaries. Finally, in Section~\ref{sec:main}, we state and prove our main result.

\section{Arithmetic of hyperelliptic curve with genus \texorpdfstring{$g=2$}{g=2}}
\label{sec:HEC} 

Let $C$ be the hyperelliptic curve defined by \eqref{def:hec} with
$$
f(X) = X^5 + b_1X^4 + b_2X^3 + b_3X^2 + b_4X + b_5 \in \Fq[X],
$$
for the finite field $\Fq$ with characteristic $\Char(\Fq) \neq 2$. Let $\Fqb$ be the algebraic closure of $\Fq$.

\subsection{Grant representation}
The Jacobian $J_C$ is an abelian variety of dimension $2$  \cite[Theorem~A.8.1.1]{hindryDiophantineGeometryIntroduction2000}. In \cite{grant1990formal}, Grant provides an embedding of $J_C$ into the projective space $\mathbb{P}^8$.
Namely, the Jacobian $J_C$ can be identified with the vanishing locus of $13$ homogenous polynomials,
\begin{equation*}
    J_C \cong V(f_2^h,\dots,f_{14}^h) = \{ z \in \mathbb{P}^8 : f_i^h(z) = 0, 2 \leq i \leq 14 \} ,
\end{equation*}
where $ f_i^h \in \Fq[\X_0, \X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X]$ are homogenized with respect to the variable $\X_0$.
For the expressions $ f_i^h$, see \cref{sec:def_eq_J}.

Let 
\begin{equation}\label{eq:Theta}
\Theta(\Fq)=\{D\in J_C(\Fq): \psi(D)=P-\cO, P\in C(\Fq) \}
\end{equation}
be the preimage of $C(\Fq) - \cO$ under $\psi$. Also write $\Theta=\Theta(\Fqb)$.
Let $\iota : J_C \rightarrow \mathbb{P}^8$ be the embedding, then
\begin{equation}\label{eq:iota}
    \iota(D) = 
        \begin{cases}
            (1:z_{11}:z_{12}:z_{22}:z_{111}:z_{112}:z_{122}:z_{222}:z)  & \text{if~} D \in J_C\setminus \Theta, \\ 
            (0:0:0:0:1:0:0:0:0) & \text{if~} D = \cO ,\\
            (0:0:0:0:-x^3: - x^2 : -x : 1 : -y) & \text{if~} D \in \Theta \setminus\cO.
            
        \end{cases}
\end{equation}

For $D\in J_C(\Fq)\setminus \Theta(\Fq)$, the components $z_{ij}, z_{ijk}$ of $\iota(D)$ can be expressed as rational functions in the coordinates $(x_1,x_2,y_1,y_2)$ of $\psi(D) = (x_1,y_1) + (x_2,y_2) - 2\cO$.
Moreover, the Mumford representation
$\eta(D)= [u,v] = [X^2 + u_1X + u_0, v_1X + v_0]$ can also be expressed with respect to $(x_1,x_2,y_1,y_2)$.
Namely, write $\psi(D) =(x_1,y_1) + (x_2,y_2) - 2\cO$ ,
if $\eta(D) = [u,v] = [X^2 + u_1X + u_0, v_1X + v_0]$, then the relation between $\iota (D)$ and $\eta(D)$ is given as follows: 

For the case $x_1 \neq x_2 $, we have 
        \begin{align}\label{eq:repr}
          z_{12} &= -x_1x_2 =-  u_0 ,  &z_{22} &= x_1+x_2 =-u_1  , \\ \notag
          z_{122}  &= \frac{x_1y_2- x_2y_1}{x_1-x_2} =  v_0 , &z_{222}& = \frac{y_1-y_2}{x_1-x_2} = v_1,  
        \end{align}
        \begin{align*}
        	z_{11}= &\frac{(x_1 + x_2)(x_1x_2)^2 + 2b_1(x_1x_2)^2 + b_2(x_1 + x_2)x_1x_2 + 2b_3x_1x_2}{(x_1-x_2)^2}\\
        	& +\frac{b_4(x_1 + x_2) + 2b_5 - 2y_1y_2}{(x_1-x_2)^2}.
        \end{align*} 
For the case $x_1 = x_2$ and $y_1 = y_2\neq 0$, we consider the identifications on page 109 of \cite{grant1990formal} and after performing some elementary computations, we obtain:
        \begin{align*}
          z_{12} &= -x_1^2 =  - u_0, & z_{22} &= 2x_1 = -u_1  \\
            z_{122} &= -\frac{f'(x_1)}{2y_1}x_1 + y_1 = v_0 , & z_{222}& = \frac{f'(x_1)}{2y_1} = v_1,  \\
            z_{11} &= \Big( \frac{f'(x_1)}{2y_1} \Big)^2 - 6x_1^3 - 4b_1x_1^2 - 2b_2x_1 - b_3
        \end{align*}
For the case $x_1 = x_2$ and $y_1 = -y_2$, we obtain $\eta(D) = [u,v] = [1,0]$ and $\iota(D) = (0:0:0:0:1:0:0:0:0)$.

The coordinates $z_{111},z_{112}$ and $z$ are given by the defining polynomials $f_4, f_3$ and $f_2$ respectively. See \cref{app:rational function}.

We denote the affine part of $J_C$ under $\iota$ with respect to $\X_0$ by $U$. Then 
$$
U = J_C\setminus \Theta.
$$
%Let $$\Fq[\mathbf{\X}] = \Fq[\X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X] $$
%
Moreover, 
by \cite[Corollary~2.15]{grant1990formal}, we have
$$
\U=V(f_2,\dots, f_7), f_i \in
\Fq[\mathbf{\X}] = \Fq[\X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X]
.
$$
Since $ J_C$ 
is a variety, it is irreducible. It follows that $ U $
is irreducible and dense, see \cite[Example 1.1.3 ]{hartshorneAlgebraicGeometry1977}. Since $J_C$ has dimension $2$, the dimension of $ U $ is also $2$.

Let $ \Fq(U) $ be the function field of $ \U $. Since $ \U $ is a dense open subset of $J_C$, one can show that $ \Fq(U) = \Fq(J_C)$, see \cite[Theorem~3.4]{hartshorneAlgebraicGeometry1977}.

Let $ h \in \Fq(U) $ be a rational function. Since $h$ is an equivalence class, we can choose a representative element $\frac{h_1}{h_2}$, where $\deg h_1$ is minimal and we define
$$
\deg h =\max\{\deg h_1, \deg h_2\}.
$$

\subsection{Group law}
We describe the group law $(Q,R) \mapsto Q+R$ for the most common case, that is, if all elements $Q,R,Q+R,Q-R$ belong to $\U$, see  \cite[Theorem 3.3]{grant1990formal}.

\begin{lemma}\label{lemma:addition_grant}
Assume that $Q,R,Q+R,Q-R \in U$.
We write
  \begin{equation}\label{eq:q}
  q(Q,R)=z_{11}(Q)-z_{11}(R)+z_{12}(Q)z_{22}(R)-z_{12}(R)z_{22}(Q).
  \end{equation}
Then there are rational functions $q_1,q_2,q_{11},q_{12},q_{22},q_{111} $ on $U\times U$ such that for $1\leq i\leq j\leq 2$ we have
\begin{equation}\label{eq:add1}
       z_{ij}(Q+R) = -z_{ij}(Q)-z_{ij}(R) + \frac{1}{4}\left( \frac{q_{i}(Q,R)}{q(Q,R)} \right) \left( \frac{q_{j}(Q,R)}{q(Q,R)} \right) -\frac{1}{4} \left( \frac{q_{ij}(Q,R)}{q(Q,R)} \right)
\end{equation} 
 and
 \begin{equation}\label{eq:add2}
    \begin{split}
         z_{111}(Q+R)= &-\frac{1}{2}z_{111}(Q)-\frac{1}{2}z_{111}(R) + \frac{3}{16}\frac{q_{1}(Q,R)q_{11}(Q,R)}{q(Q,R)^2}-\frac{1}{16}\frac{q_{111}(Q,R)}{q(Q,R)} \\
     &- \frac{1}{8} \left( \frac{q_{1}(Q,R)}{q(Q,R)} \right) ^3 
     + \frac{3}{4}(z_{11}(Q)+z_{11}(R))\frac{q_{1}(Q,R)}{q(Q,R)}, \\
    z(Q+R) = &\frac{1}{2} \left( z_{11}(Q+R)z_{22}(Q+R) - z_{12}^2(Q+R) + b_2z_{12}(Q+R) -b_4 \right).
    \end{split}
\end{equation}
\end{lemma}
 The definitions of the rational functions $q, q_1,q_2,q_{11},q_{12},q_{22},q_{111} $ are in \cref{sec:addition_formulas}. We also computed formulas for $z_{112}, z_{122}, z_{222}$ which are listed in \cref{sec:addition_formulas}.

For a fixed $ R \in \U$, we define 
 \begin{equation}\label{eq:add_fn}
 	z_{ij}^R(Q) = z_{ij}(Q+R),
 	\quad
 	z_{ijk}^R(Q) = z_{ijk}(Q+R),
 	\quad
 	z^R = z(Q+R).
 \end{equation}
If we consider $z_{ij}^R(Q), z_{ijk}^R(Q)$ to be polynomials in the variables $z_{ij}(Q), z_{ijk}(Q)$ and $z(Q)$, then it follows from \eqref{eq:add1}, \eqref{eq:add2} and \cref{sec:addition_formulas} that 
 \begin{equation} \label{eq:add_deg}
 	\deg z_{ij}^R \leq 3 \text{~  ,  ~} \deg z_{ijk}^R \leq 4 \text{~  and  ~} \deg z^R \leq 6.
 \end{equation}

 \begin{lemma}\label{lem:1}
  Let $q(Q,R)$ be defined by \eqref{eq:q} and set $q_R(Q) = q(Q,R)$.  Then for any fixed $R\in \U$,  the zero set $\{ q_R(Q)=q(Q,R)=0 \}$ has dimension one and $\Theta \pm R\subset\{q_R=0\}$. Moreover
  if $R\neq \pm R'$, then 
  \begin{equation}\label{eq:cardinality_zeros_q_1}
    |\{q_{R} = 0\} \cap \{q_{R'} = 0\} \cap \U| \leq 20.
\end{equation}
 \end{lemma}
  \begin{proof}
 Let $\bar{C}$ be the lifted curve of $C$ defined over some number field $\mathbb{K}$ and let $\mathfrak{p}$ be a prime ideal such that $C\equiv \bar{C} \bmod \mathfrak{p}$. For $\bar{C}$, the zero set of $q_{\bar{R}}$ is $\Theta(\bar{\mathbb{K}}) \pm \bar{R}$, see \cite{grant1990formal}. Thus, the points in $\Theta \pm R$ are zeros of $q_R$.
We show that 
 \begin{equation}\label{eq:subvariety}
 \{q_R=0\} \cap \U     
 \end{equation}
 has dimension one. As $\U$ is irreducible, it is enough to show that \eqref{eq:subvariety} is a proper subset. Indeed, suppose that $q_R$ vanishes on $U$. Let $Q_1,Q_2\in U(\Fq)$ corresponding to the pairs of points $(P_1,P_2)$ and $(P_1, -P_2)$ with
 \begin{equation}\label{eq:yneq0}
 P_i=(x_i,y_i),\quad  y_i\neq 0,\quad  i=1,2. 
 \end{equation}
 Then their first component in the Mumford representation are the same, but with different second components.
 
 Substituting to $q_R$, we get by the assumption that
 $$
 q_R(Q_1)=q_R(Q_2)=0
 $$
 which yields by \eqref{eq:repr} and \eqref{eq:q} that $y_1y_2=-y_1y_2$ which contradicts \eqref{eq:yneq0}. 
 
 In order to show \eqref{eq:cardinality_zeros_q_1}
 for $R\neq \pm R'$, $R,R'\in \U$, consider $\U$ as embedded in $\mathbb{A}^8$ with respect to the coordinates $z_{11},z_{12},z_{22},z_{111},z_{112},z_{122},z_{222}, z $. Then for all $R\in \U$, $q_R=0$ defines a seven dimensional affine hyperspace in $\mathbb{A}^8$. Thus the hyperspaces corresponding to $R$ and $R'$ are the same if
 $$
 z_{11}(R)=z_{11}(R'), \quad  z_{12}(R)=z_{12}(R') \quad \text{and} \quad  z_{22}(R)=z_{22}(R') 
 $$
 and thus $R = \pm R'$.

Consider the solutions
 $$
 q_R(z_{11},z_{12}, z_{22})=q_{R'}(z_{11},z_{12}, z_{22})=0, \quad $$
 where $(z_{11},z_{12}, z_{22},z_{111} ,z_{112},z_{122},z_{222}, z)\in \U$.
 As $R\neq \pm R'$, the hyperplanes defined by  $q_R$ and $q_{R'}$ are distinct. Moreover, the linear equation system
 \begin{align*}
 q_R(z_{11},z_{12}, z_{22})= z_{11}-z_{11}(R)+z_{12}z_{22}(R)-z_{12}(R)z_{22}&=0,\\
 q_{R'}(z_{11},z_{12}, z_{22})= z_{11}-z_{11}(R')+z_{12}z_{22}(R')-z_{12}(R')z_{22}&=0
 \end{align*}
 has full rank and can be transformed to the form
 \begin{equation}\label{eq:LES}
  \begin{pmatrix}
  0 & a_{11} & a_{12}\\
  1 & a_{21} & a_{22}
  \end{pmatrix}
  \begin{pmatrix}
  z_{11}\\ z_{12}\\ z_{22}
  \end{pmatrix}
  =
  \begin{pmatrix}
  c_1\\ c_2
  \end{pmatrix}
 \quad \text{with} \quad
 (a_{11},a_{12})
 %\det \begin{pmatrix}
 %  a_{11} & a_{12}\\
 %  a_{21} & a_{22}
 % \end{pmatrix}
  \neq (0,0).
 \end{equation}
 As $(z_{11},z_{12}, z_{22},z_{111} ,z_{112},z_{122},z_{222}, z)\in \U$, we can write
 \begin{align}\label{eq:z11}
     z_{11}= \frac{z_{22}z_{12}^2+2b_1 z_{12}^2-b_2z_{22}z_{12}-2b_3z_{12}+b_4z_{22}+2b_5-2y_1y_2 }{z_{22}^2+4z_{12}},
 \end{align}
 where $y_1, y_2$ are the $y$-coordinate of the points $(x_1,y_1),(x_2,y_2)\in C$ such that
 \begin{equation}\label{eq:z}
x_1x_2=-z_{12} \quad \text{and} \quad  x_1+x_2=z_{22}.
\end{equation}
 Substituting \eqref{eq:LES} into \eqref{eq:z11}, we obtain
 \begin{align*}
     2y_1y_2=&
     (z_{22}^2+4z_{12})\left(a_{21}z_{12} +a_{22}z_{22}-c_2\right)
     +z_{22}z_{12}^2+2b_1 z_{12}^2-b_2z_{22}z_{12}\\
     &-2b_3z_{12}
     +b_4z_{22}+2b_5.
 \end{align*}
Taking the square and substituting $y_i^2=f(x_i)$, ($i=1,2$), we get

\begin{align}\label{eq:sym}
     4 f(x_1)f(x_2)=
     &\Big((z_{22}^2+4z_{12})\left(a_{21}z_{12} +a_{22}z_{22}-c_2\right)
     +z_{22}z_{12}^2+2b_1 z_{12}^2\\ \notag
     &-b_2z_{22}z_{12}
     -2b_3z_{12}+b_4z_{22}+2b_5\Big)^2.
 \end{align}
 
Assume, that $a_{11}\neq 0$. The case $a_{11}=0$ and $a_{12}\neq 0$ can be handled similarly and we leave it to the reader.
 
Write
$$
g(z_{12},z_{22})= z_{22}^2 \left(a_{21}z_{12} +a_{22}z_{22}\right)+z_{22}z_{12}^2.
$$
 
First, we consider the case that
\begin{equation}\label{eq:g}
g(z_{12},z_{22})=g\left(\frac{c_1-a_{12}z_{22}}{a_{11}},z_{22}\right)
\end{equation}
has degree 3. As $f(x_1)f(x_2)$ is a symmetric polynomial in $x_1$ and $x_2$, by \eqref{eq:z} we can write
$$
f(x_1)f(x_2)=F(z_{12},z_{22}), \quad F\in \F_q[z_{12},z_{22}], \quad \deg F\leq 5.
$$
Hence, after the substitution $z_{12}=(c_1-a_{12}z_{22})/a_{11}$, the polynomial equation~\eqref{eq:sym} is non-trivial with degree at most 6. Thus, 
by~\eqref{eq:LES}, we obtain at most 12 possible solutions for $(z_{11},z_{12}, z_{22},z_{111} ,z_{112},z_{122},z_{222}, z)$.
 
Now consider the case, when~\eqref{eq:g} has degree at most two. By \eqref{eq:z} and \eqref{eq:x2} we have
\begin{equation}\label{eq:x2}
x_2=\frac{c_1-a_{12}x_1}{a_{12}-a_{11}x_1}.
\end{equation}
Then substituting it into~\eqref{eq:sym}, after clearing the denominator, we get that the left-hand side has degree 10 while the right-hand side has degree at most 9. Thus, there are at most $10$ solutions for $x_1$, and therefore at most $20$ solutions for $$(z_{11},z_{12}, z_{22},z_{111} ,z_{112},z_{122},z_{222}, z)$$ by \eqref{eq:LES}, \eqref{eq:z}  and \eqref{eq:x2} which proves the result.
\end{proof}

\begin{lemma}\label{lemma:intersection}
For $D \in U(\Fq)$ we have
$$
|\{\Theta(\Fq) + D\} \cap \Theta(\Fq)| \leq 2.
$$
\end{lemma}
\begin{proof}
We want to count the number of elements $\mathcal{P}_i \in \Theta(\Fq)$ such that 
$$
\mathcal{P}_i + D = \mathcal{P}_j \quad \text{for some~} \mathcal{P}_j \in \Theta(\Fq).
$$
Since $D \in  U(\Fq)$, we have
$$
\psi(D) = \begin{cases}
P + Q - 2\cO, &P,Q \in C(\Fq), P \neq Q, \\
2P - 2\cO, &P \in C(\Fq), \\
P' + Q' - 2\cO, &P',Q' \in C(\F_{q^2}), P', Q' \text{~are conjugates over~} \Fq. 
\end{cases}
$$
If $\psi(D) = P + Q - 2\cO, P,Q \in C(\Fq), P \neq Q$, then for 
$$
\psi(\mathcal{P}_i) = \begin{cases}
-P -\cO, &\psi(\mathcal{P}_i + D) = Q -\cO \in \Theta(\Fq) \\
-Q -\cO, &\psi(\mathcal{P}_i + D) = P -\cO \in \Theta(\Fq) ,
\end{cases}
$$
we get 2 intersection points.

If $\psi(D) = 2P - 2\cO, P \in C(\Fq)$, then for $\psi(\mathcal{P}_i) = -P-\cO$, we get $\psi(\mathcal{P}_i + D) = P -\cO \in \Theta(\Fq)$. We only get one intersection point.

If $\psi(D) = P' + Q' - 2\cO, P',Q' \in C(\F_{q^2})$ and $P', Q' \text{~are conjugates over~} \Fq$, we get no intersection points, since if $\mathcal{P}_i + D = \mathcal{P}_j, $ where $\mathcal{P}_i, \mathcal{P}_j \in \Theta(\Fq)$, with $\psi(\mathcal{P}_i) = Q_i - \cO$ and $\psi(\mathcal{P}_j) = Q_j - \cO, Q_i, Q_j \in C(\Fq)$ we get 
$$
\psi(D) = P' + Q' - 2\cO = \psi(\mathcal{P}_j + (-\mathcal{P}_i)) = Q_j - Q_i - 2\cO
$$
which is a contradiction.
\end{proof}

\begin{prop}\label{prop:main_1}
	Let $ h \in \Fq(U)$ be a rational function with pole divisor of the form $ n\Theta$, for $n \geq 1$. Let $ W_0 \in J_C(\Fq) $ and $ D \in J_C(\Fq) $ be an element of order $ t $. Let $L$ be a positive integer with 
	\begin{equation}\label{eq:prop-L}
	L < \min\left\{ \frac{t-1}{2} - |\Theta(\Fq)| , \frac{|\Theta(\Fq)| - 2}{20}\right\}. 
	\end{equation}
	Let $ j_0< j_1< \dots< j_{L} \leq L + |\Theta(\Fq)|+1$ be positive integers such that $j_iD + W_0 \not \in \Theta(\Fq)$. Let $ c_0, \dots , c_L  \in \Fq$ with $ c_L \neq 0 $. Then the rational function $H \in \Fq(U)$, with 
	\begin{equation*}
	H(Q) = \sum_{l=0}^{L} c_l h(Q + j_l D + W_0)
	\end{equation*}
	is non-constant and has degree 
	$$ 
	\deg H \leq 6(L +1) \deg h.
	$$
\end{prop}
\begin{proof}
First, we consider the case when $ W_0 = 0 $. Defining the function $ h_D : Q \mapsto h(Q + D) $ yields 
$$
H(Q) = \sum_{l=0}^{L-1} c_l h_{j_lD}(Q) + c_L h_{j_LD}(Q).
$$ 
We show that there exists $ Q \in \U $ such that it is a pole of $ h_{j_LD} $, but not a pole of any other terms $ h_{j_lD},$ for $ j_l<j_L $.

Observe that $h_{j_LD} $ has a pole at $ Q $ when $ Q \in \Theta(\Fq) - j_LD $.
From \cref{lem:1}, we know that $ \Theta(\Fq) -j_LD \subseteq \{q_{j_LD} = 0\} $.   Moreover, it follows from  \cref{lemma:intersection}, that
\begin{equation}\label{eq:card_j_L}
    |\Theta(\Fq) -j_LD|\geq |\Theta(\Fq)|-2.
\end{equation}

By \eqref{eq:prop-L}, we have
$$
   j_l + j_L \leq j_{L-1} + j_L \leq 2L + 2|\Theta(\Fq)| + 1 < t,
$$
for any $l< L$, whence $ j_lD \neq -j_LD $ for $ j_l<j_L $. 
Then by Lemma \ref{lem:1}, we obtain
$$
\left| \Big((\Theta(\Fq) - j_LD) \cap \U \Big) \cap \{q_{j_lD} = 0\} \right|\leq
\left| \{q_{j_LD}=0\} \cap \{q_{j_lD} = 0\} \cap \U\right|\leq 20 .
$$
Thus, by \eqref{eq:card_j_L} we have that
\begin{align}\label{eq:set}
		 &\left|\Big((\Theta(\Fq) - j_LD) \cap \U \Big) \setminus \left( \bigcup\limits_{l=0}^{L-1} \{q_{j_lD} = 0 \} \right)\right| \\ \notag
		 =&\left| \bigcap\limits_{l=0}^{L-1} \bigg( \Big((\Theta(\Fq) - j_LD) \cap \U \Big)  \setminus \{q_{j_lD} = 0\} \bigg)\right| \geq |\Theta(\Fq)|-2-20L.
\end{align}
By \eqref{eq:prop-L}, the set in \eqref{eq:set} is non-empty, and thus there exists a point $Q$ which is a pole of $h_{j_LD}$ but not a pole of any other term of $H$. Hence, $H$ is non-constant.

In the case when $ W_0 \neq 0 $, we can define 
$$
\widetilde{H}(Q) = H(Q-W_0) = \sum_{l=0}^{L} c_lh(Q + j_l D + W_0-W_0).
$$
Then, by the case $\widetilde{W}_0 = W_0 - W_0 = 0$, we obtain that $ \widetilde{H}$ is non-constant. 
Therefore, there exist $ Q_1,Q_2 \in \U $ such that $ \widetilde{H}(Q_1) \neq \widetilde{H}(Q_2) $. Hence, we obtain
$$
	H(Q_1-W_0) \neq H(Q_2 - W_0).
$$
This proves that $ H $ is non-constant.

To estimate the degree of $ H $, we first estimate the degree of the functions $ h_{j_lD}$. Let $l$ be arbitrary, define $R = j_l D + W_0$, 
Let $ z_{ij}^{R}, z_{ijk}^{R}, z^{R} $ be as defined in \eqref{eq:add_fn}, then we can write $ h_{R}(Q) $ as 
\begin{equation*}
	h_{R}(Q) =h\big(Q + R\big) =h\big(z_{11}^{R}(Q), \dots , z_{222}^{R}(Q),z^{R}(Q)\big).
\end{equation*}
It follows from \eqref{eq:add_deg} that
\begin{align*}
	\deg h_{R} \leq (\deg h) \big(\max \{ \deg z_{ij}^{R}, \deg z_{ijk}^{R}, \deg z^R \}\big) = 6 \deg h,
\end{align*}
and thus
\begin{equation*}
	\deg H \leq \deg \bigg( \sum_{l=0}^{L} c_l h_{R} \bigg) \leq 6(L+1) (\deg h).
\end{equation*}
In particular, the degree does not depend on $ W_0 $.

\end{proof}

\section{Further preliminaries}\label{sec:preliminaries}

\subsection{Linear complexity}

We need the following result on linear complexity, see \cite[Lemma~ 6]{langeCertainExponentialSums2005}.
\begin{lemma}\label{lemma:linComp}
Let $(s_n)$ be a linear recurrent sequence of order $L$ over any finite field $\Fq$ defined by a linear recursion
$$
s_{n+L}=c_0s_n +\dots + c_{L-1}s_{n+L-1}, \quad n\geq 0.
$$
Then for any $T\geq L+1$ and pairwise distinct positive integers $j_1,\dots, j_{T}$, there exist $a_1,\dots, a_T \in \Fq$, not all equal to zero, such that
$$
\sum_{i=1}^Ta_is_{n+j_i}=0, \quad n\geq 0.
$$
\end{lemma}

\subsection{Vanishing loci of polynomials}

Let $\K$ be a field and for polynomials $f_1,\dots, f_k\in \K[\mathbf{X}]$ with $\mathbf{X}=(X_1,\dots, X_n)$. We denote
$$
V_\K(f_1,\dots, f_k)=\{\mathbf{x}\in\K^n: f_1(\mathbf{x})=\dots=f_k(\mathbf{x})=0 \} 
$$
and
$$
V(f_1,\dots, f_k)=V_{\overline{\K}}(f_1,\dots, f_k).
$$
The following result is a multidimensional version of B\'ezout's Theorem, see
\cite[Theorem~3.1]{schmid1995affine}.

\begin{lemma}\label{lemma:bezout}
Assume, that  $f_1,\dots, f_k\in \K[\mathbf{X}]$ and $\dim V(f_1,\dots, f_k) \leq 0$. Then
$$
| V(f_1,\dots, f_k)|\leq \prod_{i=1}^k \deg f_i.
$$
\end{lemma}

\section{Main results}\label{sec:main}

Recall, that for $D\in J_C(\Fq)$ we have defined the sequence $(W_n)$ recursively by \eqref{eq:S}, namely
$$
W_n=D + W_{n-1}=nD+ W_0, \quad n=1,2,\dots,
$$
with some initial value $W_0\in J_C(\Fq)$. We can assume, that $D,W_0\in J_C(\Fq)$ and thus  all sequence elements are defined over $\Fq$. 
Let $h\in \Fq(J_C)$ and consider the sequence
\begin{equation}\label{eq:w}
w_n=h(W_n), \quad n=0,1,\dots
\end{equation}
with the convention that, if $h$ is not defined at $W_n$, we set $w_n=0$. Clearly, $(W_n)$ and $(w_n)$, are purely periodic sequences, and if $t$ is the order of $D$ in $J_C(\Fq)$, then $t$ is the period length of $(W_n)$. However, $(w_n)$ may have a smaller period length.

\begin{theorem}\label{thm:main}
	Let $ C $ be a hyperelliptic curve of genus $ 2 $. Let $ h \in \Fq(U) $ be a rational function in the function field of the Jacobian with pole divisor of the form $ n\Theta$, for $n \geq 1$.  If $W_0 \in J_C(\Fq)$ and $ D \in J_C(\Fq) $ is of order $ t $ and $ w_n$ is defined by \eqref{eq:w}, then 
	\begin{equation*}
		L(w_n,N) \geq \left\lfloor c 
		   \frac{\min\{t,N\}}{ q \deg h }\right\rfloor
	\end{equation*}
	for some absolute constant $c>0$. 
\end{theorem}

Theorem~\ref{thm:main} yields a lower bound on the linear complexity of the components \eqref{eq:coordinates} in the Mumford representation of $(W_n)$. Clearly, $u_2$ is constant on $\U$. However, applying the result for $h=-z_{22}, -z_{12}, z_{122}$ and $z_{222}$ and observing that $\Theta$ is a pole of them (cf. \eqref{eq:iota}), we get a lower bound on the linear complexity of the components $u_1,u_0, v_1$ and $v_0$ by \eqref{eq:repr}. As all functions $-z_{22}, -z_{12}, z_{122}$ and $z_{222}$ have degree one, we get \eqref{eq:example}.

The result is non-trivial if $t>c q$ for some constant which may depend on $\deg h$. However, the most important case is that $t$ is close to $|J_C(\Fq)|$ which is $(1+o(1))q^2$ by \eqref{eq:size_of_jac}.

\begin{proof}
Let $\Theta(\Fq)$ be defined as in \eqref{eq:Theta}. Then by \eqref{eq:number_of_points} we have $|\Theta(\Fq)|=|C(\Fq)|=q+O(q^{1/2})$. We can assume
\begin{equation}\label{eq:assumption2}
    N \geq 6|\Theta(\Fq)|+2q +3
\end{equation}
and
\begin{equation}\label{eq:assumption}
    L<\min\left\{\frac{t-2-5|\Theta(\Fq)|-2q}{2|\Theta(\Fq)|}, \frac{N-3-6|\Theta(\Fq)|-2q}{2|\Theta(\Fq)|+1}, \frac{t-1}{2}-|\Theta(\Fq)|, \frac{|\Theta(\Fq)|-2}{20} \right\}
\end{equation}
since otherwise, we can choose the absolute constant $c$ small enough so that the theorem holds trivially.

Let $L$ be the $N$th linear complexity of the sequence $(w_n)$ and let $c_0,\dots, c_L\in \Fq$ such that
$$
w_{n+L}=c_0w_n+\dots+c_{L-1}w_{n+L-1}, \quad 0\leq n\leq N-L-1.
$$
Let $(s_n)$ be the infinite linear recurrent sequence with $s_n=w_n$ for $0\leq n<N$ and 
$$
s_{n+L}=c_0s_n+\dots+c_{L-1}s_{n+L-1}, \quad n \geq 0.
$$
Let 
\begin{equation}\label{eq:j_L}
1\leq j_0<\dots<j_{L}\leq |\Theta(\Fq)|+L+1
\end{equation}
be the smallest integers such that $W_{j_i}\not\in\Theta, 0 \leq i \leq L $. 
Then by \cref{lemma:linComp}, there exist $a_0\dots, a_{L}\in\Fq$ not all equal to zero, such that
$$
\sum_{i=0}^La_is_{n+j_i}=0, \quad n\geq 0,
$$
whence, 
\begin{equation}\label{eq:rec}
\sum_{i=0}^{L}a_iw_{n+j_i}=0, \quad 0\leq n <  \min \left\{ N-j_L, t \right\} 
\end{equation}
as $w_n=s_n$ for $0\leq n\leq N$.

We define 
\begin{equation}\label{eq:T}
T = \min \left\{ N-j_L, t \right\}
\end{equation}
and put
$$
\mathcal{N} = \Big\{ 0 \leq n<T : \ nD,  nD \pm (j_i D + W_0)\not\in \Theta(\Fq), {~\text{for}~} 0\leq i \leq L\Big\}.
$$ 

We observe that there are at most $|\Theta(\Fq)|$ points $nD\in\Theta(\Fq)$ and similarly for each $j=\pm j_i$ ($i=0,\dots, L$), there are  $2|\Theta(\Fq)|$ elements $nD \pm (j_iD + W_0) \in\Theta(\Fq)$. Hence, by \eqref{eq:assumption} we have,
\begin{equation}\label{eq:N}
|\mathcal{N}| \geq T - 2L|\Theta(\Fq)| - 3|\Theta(\Fq)|>2|\Theta(\Fq)|+2(q+1).
\end{equation}

Define
\begin{equation}\label{min_zeros_F}
	H(Q) = \sum_{i=0}^{L} a_{i} h( Q+ j_{i}D +W_0). 
\end{equation}
For $Q=nD$, $n\in\cN$, using \cref{lemma:addition_grant}, we see that $H$ is well-defined.
By \eqref{eq:rec}, $H$ vanishes on $Q=nD$, $n\in\cN$. We give an upper bound on the number of zeros of $H$ to get the results together with \eqref{eq:T} and \eqref{eq:N}.

Let us fix a representation of $H$ as a rational function $G_1/G_2\in\Fq(\mathbf{\X})$. Then this set is finite and contains the zeros of $H$ and thus
\begin{equation}\label{eq:N_V}
    |\cN|\leq |  V_{\Fq}(f_2,\dots,f_7,G_1 )|
\end{equation}
where $f_2, \dots, f_7 $ are the defining equations of $\U$. See \cref{sec:def_eq_J}.

In order to estimate the size of \eqref{eq:N_V}, for $r\in\Fq$, write $g_r(\mathbf{\X})=\X_{12}-r$. Clearly,
\begin{equation}\label{eq:slicing}
  V_{\Fq}(f_2,\dots,f_7,G_1)\subset \bigcup_{r\in\Fq}   V(f_2,\dots,f_7,G_1, g_r).
\end{equation}
We claim that for any $r\in\Fq$,
\begin{equation}\label{eq:finite_dim}
    \dim(V(f_2,\dots,f_7,G_1,g_r)) = 0.
\end{equation}
From \cref{prop:main_1}, we know that $H$ is non-constant, therefore,
$$
\dim(V(f_2,\dots,f_7,G_1)) = 1.
$$
Hence, if \eqref{eq:finite_dim} were not true, then $\dim(V(f_2,\dots,f_7,G_1,g_r)) = 1$ and so
$$
g_r \in \langle f_2,\dots,f_7,G_1 \rangle.
$$
Let $D \in \U$ such that $G_1(D) = 0$. Then we also have $g_r(D) = 0$. As all $nD$ for $n\in\cN$ are zeros of $G_1$, it has at least 
\begin{equation}\label{eq:claim}
    2|\Theta(\Fq)|+2(q+1) + 1
\end{equation}
zeros by \eqref{eq:N}.
On the other hand, if $D=[P+Q-2\cO]\in U(\Fq)$, with $P=(x_P,y_P), Q=(x_Q,y_Q)$, is a zero of $g_r$, then
$$
g_r(D)= z_{12}(D)-r=-x_Px_Q-r=0.
$$
If $P,Q\in C(\Fq)$, then $P$ determines $Q$ apart from sign, therefore there are at most $2|\Theta(\Fq)|$ such zeros. On the other hand, if $P,Q\in C(\F_{q^2})$ and $P$ and $Q$ are conjugated, then we must have $x_Q=x_P^q$ and thus $x_P^{q+1}=-r$. As there are at most $q+1$ solutions, there are at most $2(q+1)$ such divisors. Then it yields that $g_r$ has at most $2|\Theta(\Fq)|+2(q+1)$ zeros over $\Fq$, which contradicts \eqref{eq:claim}. Then we have proved \eqref{eq:finite_dim}.

By Lemma~\ref{lemma:bezout} and \eqref{eq:finite_dim} we have
$$
|V(f_2,\dots,f_7,G_1,g_r)|\leq \prod_{i=2}^7 \deg f_i \cdot \deg G_1\leq 216 \deg G_1, 
$$
and thus by \eqref{eq:slicing},
\begin{equation}
       V_{\Fq}(f_2,\dots,f_7,G_1)\leq  216 q\deg G_1\leq 216 q\deg H. 
\end{equation}
Then it follows from Proposition~\ref{prop:main_1}, \eqref{eq:N} and \eqref{eq:N_V} that
\begin{align*}
T-2L|\Theta(\Fq)|-3|\Theta(\Fq)| \leq |\cN| \leq  216 q \deg H \leq 1296(L+1)q \deg h. 
\end{align*}
Whence \eqref{eq:j_L} and \eqref{eq:T} yields the result with some absolute constant $c>0$.
\end{proof}

\section{Comments}
The usage of elliptic curves in pseudorandom number generation is an extensive research direction,
see for example the survey paper \cite{survey_Shparlinski} for a discussion of properties of pseudorandomness of the elliptic curve case of the sequence \eqref{eq:S} and further references as well as other constructions.

Despite some results on the application of hyperelliptic curves in pseudorandom number generation (see \cite{ShparlinskiLange05,Rezaeian,Reza2}) this
line of research has never been studied systematically. In Theorem~\ref{thm:main}, we obtain lower bounds on the linear complexity of the sequence derived from~\eqref{eq:S}. 

The bound we obtain is non-trivial and sufficient for application, however we conjecture that stronger bound for the $N$-th linear complexity holds. Numerical experiments suggest that if $h$ is a coordinate function in \eqref{eq:coordinates}, then for most cases $L(w_n,t)=\lceil t/2 \rceil$.

Additional to the $N$-th linear complexity,
other pseudorandom properties, like uniform distribution, also need to be investigated. The main tool would be the higher genus analogue of \cite{KohelShparlinski} where the authors obtained bounds on exponential sums over elliptic curves.

 \section*{Acknowledgement}
The authors wish to thank Arne Winterhof for the valuable discussions and Igor Shparlinski for useful comments.
The authors were supported by the Austrian Science Fund
Project P31762.

\appendix
\section{}
\subsection{Defining equations of the Jacobian}
\label{sec:def_eq_J}
Let 
$$
S= \K[\X_0, \X_{11},\X_{12},\X_{22},\X_{111},\X_{112},\X_{122},\X_{222},\X] $$ 
be a polynomial ring over field $\K$, with characteristic $p \neq 2$. Following \cite{grant1990formal}, in particular Theorem 2.5, Theorem 2.11 and Corollary 2.15, we define $f_i $ as follows: 
\begin{align*}
    f_1 = &\X^2 + \X_{11}^2\X_{12} + b_1\X_{11}^2\X_{22} +   b_2\X_{11}^2\X_{12}\X_{22} - b_3\X_{11}\X_{22}^2 + b_4\X_{12}\X_{22}^2 \\
    &-b_5\X_{22}^3 + 2b_1\X\X_{11} -2b_2\X\X_{12} + 2b_3\X\X_{22}+ (b_3-b_1b_2)\X_{11}\X_{12} \\
    &+(b_2^2-b_1b_3)\X_{11}\X_{22} + (b_1b_4-b_2b_3-b_5)\X_{12}\X_{22}-b_1b_5\X_{22}^2 \\
    & +2(b_1b_3 -b_2^2)\X + (b_1b_4-b_5)\X_{11} +b_2(b_2^2 - b_1b_3 )\X_{12} \\
    &(b_3b_4 - b_2b_5)\X_{22} + b_1b_3b_4 - b_2^2b_4 - b_3b_5, \\
    f_2 =&2\X - \X_{11}\X_{22} + \X_{12}^2 - b_2\X_{12} + b_4, \\
    f_3 =&\X_{112} -\X_{222}\X_{12} + \X_{122}\X_{22}, \\
    f_4 =&\X_{111} + \X_{222}\X_{11} + \X_{122}\X_{12} - 2\X_{112}\X_{22} - 2b_1\X_{112} + b_2\X_{122}, \\
    f_5 =&\X_{122}^2 - \X_{11}\X_{22}^2 + 2\X\X_{22} + \X_{11}\X_{12} - b_1\X_{11}\X_{22} - b_2\X_{12}\X_{22} \\
    & +2b_1\X - b_1b_2\X_{12} + b_4\X_{22} + b_1b_4 - b_5 , \\
    f_6 =& \X_{222}^2 - \X_{22}^3 - \X_{12}\X_{22} - b_1\X_{22}^2 - \X_{11} - b_2\X_{22} - b_3 , \\
    f_7 =&\X_{122}\X_{222} - \X_{12}\X_{22}^2 + \X -b_2\X_{12} - b_1\X_{12}\X_{22}, \\
    f_8 =&\X_{111}^2 - \X_{11}^3 - b_3\X_{11}^2 -b_4\X_{11}\X_{12} + 3b_5\X_{11}\X_{22} + 2b_5\X \\
    &+(4b_1b_5 - b_2b_4)\X_{11} - 3b_2b_5\X_{12} + (4b_3b_5-b_4^2)\X_{22} \\
    &4b_1b_3b_5 + b_4b_5 - b_1b_4^2 - b_2^2b_5, \\
    f_9 =& -\X_{111}\X_{112} + b_1\X_{111}\X_{122} -b_2\X_{112}\X_{122} + b_3\X_{112}\X_{222} \\
    &-b_4\X_{122}\X_{222} + b_5\X_{222}^2 -\X^2 -b_1 \X \X_{11} + b_2 \X \X_{12} - b_3 \X \X_{22} \\
    &-b_3\X_{11}\X_{12} + b_1b_3\X_{11}\X_{22} - (b_5 + b_1b_4)\X_{12}\X_{22} + 2b_1b_5\X_{22}^2 \\
    & -2(b_1b_3 + b_4)\X  + (2b_2b_4 + b_1b_2b_3 +b_1b_5 - b_3^2 -b_1^2b_4)\X_{12} \\
    &- 2b_5\X_{11} + 2b_5(b_1^2 - b_2)\X_{22} + b_1b_2b_5 - b_1b_3b_4 - 2b_3b_5 , \\
    f_{10} =& \X_{122}^2 - \X_{111}\X_{122} + \X_{11}\X - b_3\X_{11}\X_{22} + 2b_4\X_{12}\X_{22} -3b_5\X_{22}^2 \\
    &+2b_3\X + (b_1b_4 - b_2b_3 - b_5)\X_{12} - 2b_1 b_5 \X_{22} + b_3b_4 - b_2b_5, \\
    f_{11} =&\X_{111}\X_{222} - \X_{112}\X_{122} -2\X\X_{12} + \X_{11}^2 - 2b_1 \X_{11}\X_{12} \\
    &+3b_2 \X_{11}\X_{22} - 2b_3 \X_{12}\X_{22} + b_4\X_{22}^2 -5b_2\X +b_3\X_{11} \\
    &+ (3b_2^2 - 2b_1b_3)\X_{12} + (b_1b_4 - b_5)\X_{22} - 2b_2b_4, \\
    f_{12}=&\X_{122}^2 - \X_{112}\X_{222} + \X_{22}\X + 2\X_{11}\X_{12} - b_1\X_{11}\X_{22} + 2b_1\X \\
    &+(b_3-b_1b_2)\X_{12} + b_1b_4 - b_5, \\
    f_{13} =&\X_{111}\X_{12} -\X_{112}\X_{11} - b_4\X_{122} + 2b_5\X_{222}, \\
    f_{14} =& 2\X_{122}\X_{11} - \X_{112}\X_{12} -\X_{111}\X_{22} - b_2\X_{112} + 2b_3\X_{122} - b_4\X_{222}.
\end{align*}
One can show that $f_1\in \langle f_5,f_6,f_7\rangle$
and the vanishing locus of these polynomials homogenized with respect to the variable $\X_0$ forms a set of defining equations for the Jacobian $J_C$, i.e 
$$
J_C = V(f_2^h,\dots,f_{14}^h) = \{ z \in \mathbb{P}^8(\bar{\K}) : f_i^h(z) = 0, 2 \leq i \leq 14 \}
$$

\subsection{Rational embedding for the Jacobian} \label{app:rational function}
Recall \cref{eq:iota}, where for any $D \in U$ with $ \psi(D) = (x_1,y_1)+ (x_2,y_2)- 2\cO $, the image of $D$ under the embedding $\iota : J_C \rightarrow \mathbb{P}^8$ is given by
\begin{equation*}
    \iota(D) = (1:z_{11}:z_{12}:z_{22}:z_{111}:z_{112}:z_{122}:z_{222}:z)
\end{equation*}
where, $z_{ij}, z_{ijk}$ are rational functions in the coordinates $x_1, x_2, y_1, y_2$ of $\psi(D)$ as shown below. For more details, see \cite[Equation 1.4]{grant1990formal}.  

\begin{align*}
    z_{11}= &\frac{(x_1 + x_2)(x_1x_2)^2 + 2b_1(x_1x_2)^2 + b_2(x_1 + x_2)x_1x_2 + 2b_3x_1x_2}{(x_1-x_2)^2}\\
    & +\frac{b_4(x_1 + x_2) + 2b_5 - 2y_1y_2}{(x_1-x_2)^2},\\
    z_{12} =& -x_1x_2,\\
    z_{22} =& x_1 + x_2, \\
    z_{111} =& \frac{y_2 \Psi(x_1,x_2) - y_1 \Psi(x_2,x_1)}{(x_1 - x_2)^3}, \text{~where,}\\
    \Psi(x_1,x_2) = &4b_5 + b_4(3x_1 + x_2) + 2b_3x_1(x_1 + x_2) + b_2x_1^2(x_1 + 3x_2) \\
    & + 4b_1x_1^3x_2 + x_1^3x_2(3x_1 + x_2),\\
    z_{112} =& \frac{y_1x_2^2 - y_2x_1^2}{x_1- x_2}, \\
    z_{122} =& -\frac{y_1x_2 - y_2x_1}{x_1- x_2}, \\
    z_{222} =& \frac{y_1 - y_2}{x_1- x_2}. \\
    z =& \frac{1}{2}(z_{11}z_{22} - z_{11}^2 + b_2z_{12} - b_4 )
\end{align*} 
We consider $U$ embedded in $\mathbb{A}^8$ by considering the isomorphism given by 
\begin{equation*}
    (1:z_{11}:z_{12}:z_{22}:z_{111}:z_{112}:z_{122}:z_{222}:z) \mapsto (z_{11},z_{12},z_{22},z_{111},z_{112},z_{122},z_{222},z)
\end{equation*}

\subsection{Addition formulas}
\label{sec:addition_formulas}
In \cref{lemma:addition_grant}, we stated the group law on $U$ as given in \cite[Theorem 3.3]{grant1990formal}. This reference gives instructions as to how one could compute $z_{112}(Q+R),z_{122}(Q+R),z_{222}(Q+R)$ but does not state them explicitly in the paper. To estimate the degree of these functions, we computed the formulas for $z_{112}(Q+R),z_{122}(Q+R),z_{222}(Q+R)$ as follows:  
\begin{align*}
     z_{112}(Q+R)= &-\frac{1}{2}z_{112}(Q)-\frac{1}{2}z_{112}(R) + \frac{1}{16}\frac{q_{2}(Q,R)q_{11}(Q,R)}{q(Q,R)^2}\\ &+\frac{1}{8}\frac{q_{1}(Q,R)q_{12}(Q,R)}{q(Q,R)^2} -\frac{1}{16}\frac{q_{112}(Q,R)}{q(Q,R)}
     - \frac{1}{8}\frac{q_{2}(Q,R)(q_{1}(Q,R))^2}{q(Q,R)^3} \\  
     &+ \frac{3}{8}(z_{11}(Q)+z_{11}(R))\frac{q_{2}(Q,R)}{q(Q,R)} + \frac{3}{8}(z_{12}(Q)+z_{12}(R))\frac{q_{1}(Q,R)}{q(Q,R)} \\
     z_{122}(Q+R)= &-\frac{1}{2}z_{122}(Q)-\frac{1}{2}z_{122}(R) + \frac{1}{16}\frac{q_{1}(Q,R)q_{22}(Q,R)}{q(Q,R)^2}\\ &+\frac{1}{8}\frac{q_{2}(Q,R)q_{12}(Q,R)}{q(Q,R)^2} -\frac{1}{16}\frac{q_{122}(Q,R)}{q(Q,R)} \\
     &- \frac{1}{8}\frac{q_{1}(Q,R)(q_{2}(Q,R))^2}{q(Q,R)^3}  
     + \frac{3}{4}(z_{12}(Q)+z_{12}(R))\frac{q_{2}(Q,R)}{q(Q,R)} \\
     z_{222}(Q+R)= &-\frac{1}{2}z_{222}(Q)-\frac{1}{2}z_{222}(R) + \frac{3}{16}\frac{q_{2}(Q,R)q_{22}(Q,R)}{q(Q,R)^2}-\frac{1}{16}\frac{q_{222}(Q,R)}{q(Q,R)} \\
     &- \frac{1}{8} \left( \frac{q_{2}(Q,R)}{q(Q,R)} \right) ^3 
     + \frac{3}{4}(z_{22}(Q)+z_{22}(R))\frac{q_{2}(Q,R)}{q(Q,R)} \\
     z(Q+R) =& \frac{1}{2}(z_{11}(Q+R)z_{22}(Q+R) - z_{11}^2(Q+R) + b_2z_{12}(Q+R) - b_4 )
\end{align*}

To evaluate the addition formulas from \cref{lemma:addition_grant}, we need the following functions:
\newcommand{\p}[2]{z_{#1#2}}
\newcommand{\pp}[3]{2z_{#1#2#3}}
\newcommand{\po}{(2z - b_2 z_{12} + b_4)}

\begin{align*}
    q(Q,R)= &z_{11}(Q)-z_{11}(R)+z_{12}(Q)z_{22}(R)-z_{12}(R)z_{22}(Q), \\
    q_1(Q,R) = &\pp111(Q) - \pp111(R) + \pp112(Q)\p22(R) - \pp112(R)\p22(Q) \\
    & +\pp122(R)\p12(Q) - \pp122(Q)\p12(R), \\ 
    q_2(Q,R) =& \pp112(Q) - \pp112(R) + \pp122(Q)\p22(R) - \pp122(R)\p22(Q) \\
    & +\pp222(R)\p12(Q) - \pp222(Q)\p12(R), \\
    q_{11}(Q,R)=& 4b_3 q(Q,R) + 4b_4(\p12(Q)-\p12(R)) + 4(\po(Q)\p12(R))\\
    &-4(\po(R)\p12(Q))-8b_5(\p22(Q) - \p22(R)) \\ &+2(\pp112(Q)\pp122(R) - \pp112(R)\pp122(Q)),\\
    q_{12}(Q,R) =& 4b_3(\p12(Q)-\p12(R)) + 2b_2(\p12(Q)\p22(R)) \\ &-2b_2(\p12(R)\p22(Q)) - 4(\p11(Q)\p12(R)-\p11(R)\p12(Q))\\
    & +2(\po(Q)\p22(R)-\po(R)\p22(Q)) \\
    & - 2b_4(\p22(Q) - \p22(R)) + \pp222(R)\pp112(Q)-\pp222(Q)\pp112(R), \\
    q_{22}(Q,R) =& 8b_1(\p12(Q)\p22(R)-\p12(R)\p22(Q)) + 4b_2\p12(Q)\\
    & -4b_2\p12(R) -8(\p11(Q)\p22(R)-\p11(R)\p22(Q)) \\
    & - 4(\po(Q)-\po(R)) \\
    & + 2(\pp122(Q)\pp222(R) -\pp122(R)\pp222(Q)),\\
    q_{111}(Q,R) =& 4b_3q_1(Q,R) \\
    & + 4(\pp111(Q)\p22(Q)\p12(R)-\pp111(R)\p22(R)\p12(Q)) \\
    & + \pp122(R)(2\p12(Q)(6\p11(Q) - 2\p11(R)+4b_3)-4b_4\p22(Q))\\
    & - \pp122(Q)(2\p12(R)(6\p11(R) - 2\p11(Q)+4b_3)-4b_4\p22(R))\\
    & + \pp112(Q)(\p12(R)(12\p12(R) - 8\p12(Q)+4b_2)+4b_4)\\
    & - \pp112(R)(\p12(Q)(12\p12(Q) - 8\p12(R)+4b_2)+4b_4)\\
\end{align*}
\subsubsection{Formulas for $q_{ijk}$}
\label{sec:addition_formulas_extended}
    There are multiple ways to compute formulas for $q_{ijk}$, e.g. 
    $$
    q_{ijk} (Q,R) = \mathcal D_i(q_{jk})(Q,R) = \mathcal D_j(q_{ik})(Q,R) = \mathcal D_k(q_{ij})(Q,R)
    $$
    where $\mathcal D_i$ is the differential operator as defined in the proof of \cite[ Theorem 3.3 ]{grant1990formal}. We used the Python package \texttt{SymPy} to compute the formulas and chose the following expressions for $q_{ijk}$:
	\begin{align*}
		q_{111} &= \mathcal D_1(q_{11}), \\
		q_{112} &= \mathcal D_2(q_{11}), \\
		q_{122} &= \mathcal D_2(q_{12}), \\
		q_{222} &= \mathcal D_2(q_{22}).
	\end{align*}

	\begin{align*}
	q_{111}(Q,R) = &~\pp112(R)
	\left(- 12 \p12(Q)^{2} + \p12(Q) \left(8 \p12(R) - 4 b_{2}\right) - 4 \p22(Q) b_{3} - 4 b_{4}
	\right) \\ 
	&+\pp111(R)
	\left(- 4 \p12(Q) \p22(R) - 4 b_{3}
	\right) \\ 
	&+\pp111(Q)
	\left(4 \p12(R) \p22(Q) + 4 b3
	\right) \\ 
	&+\pp112(Q)
	\left(- 8 \p12(Q) \p12(R) + 12 \p12(R)^{2} + 4 \p12(R) b_{2} + 4 \p22(R) b_{3} + 4 b_{4}
	\right) \\ 
	&+\pp122(Q)
	\left(4 \p11(Q) \p12(R) - 12 \p11(R) \p12(R) - 12 \p12(R) b_{3} + 4 \p22(R) b_{4}
	\right) \\ 
	&+\pp122(R)
	\left(12 \p11(Q) \p12(Q) + \p12(Q) \left(- 4 \p11(R) + 12 b_{3}\right) - 4 \p22(Q) b_{4}
	\right) \\
	q_{112}(Q,R) = &~\pp222(Q)
	\left(4 \p11(Q) \p12(R) - 4 \p12(R) b_{3} - 8 b_{5}
	\right) \\ 
	&+\pp112(Q)
	\left(- 4 \p11(R) + 4 \p12(R) \p22(Q) + \p12(R) \left(12 \p22(R) + 8 b_{1}\right) + 4 b_{3}
	\right) \\ 
	&+\pp112(R)
	\left(4 \p11(Q) + \p12(Q) \left(- 12 \p22(Q) - 4 \p22(R) - 8 b_{1}\right) - 4 b_{3}
	\right) \\ 
	&+\pp122(Q)
	(- 8 \p11(R) \p22(R) - 8 \p12(Q) \p12(R) - 4 \p12(R)^{2} - 4 \p12(R) b_{2} \\
	&+ 4 \p22(R) b_{3} + 4 b_{4} ) \\ 
	&+\pp122(R)
	\left(8 \p11(Q) \p22(Q) + 4 \p12(Q)^{2} + \p12(Q) \left(8 \p12(R) + 4 b_{2}\right) \right.\\
    &- \left. 4 \p22(Q) b_{3} - 4 b_{4} \right) \\ 
	&+\pp222(R)
	\left(\p12(Q) \left(- 4 \p11(R) + 4 b_{3}\right) + 8 b_{5}
	\right) \\
	q_{122}(Q,R) = &~\pp112(R)
	\left(- 6 \p22(Q)^{2} + \p22(Q) \left(- 2 \p22(R) - 4 b_{1}\right) - 2 b_{2}
	\right) \\ 
	&+\pp122(R)
	\left(- 4 \p11(Q) + \p22(Q) \left(4 \p12(R) - 2 b_{2}\right) - 4 b_{3}
	\right) \\ 
	&+\pp222(Q)
	\left(2 \p11(Q) \p22(R) - 4 \p11(R) \p22(R) - 2 \p12(R)^{2} - 4 \p12(R) b_{2} - 2 b_{4}
	\right) \\ 
	&+\pp112(Q)
	\left(2 \p22(Q) \p22(R) + 6 \p22(R)^{2} + 4 \p22(R) b_{1} + 2 b_{2}
	\right) \\ 
	&+\pp222(R)
	\left(4 \p11(Q) \p22(Q) - 2 \p11(R) \p22(Q) + 2 \p12(Q)^{2} + 4 \p12(Q) b_{2} + 2 b_{4}
	\right) \\ 
	&+\pp122(Q)
	\left(4 \p11(R) - 4 \p12(Q) \p22(R) + 2 \p22(R) b_{2} + 4 b_{3}
	\right) \\
	q_{222}(Q,R) = &~\pp222(R)
	\left(- 12 \p11(Q) + 4 \p11(R) + \p12(Q) \left(12 \p22(Q) + 16 b_{1}\right)
	\right) \\ 
	&+\pp122(R)
	\left(- 8 \p12(Q) - 8 \p12(R) - 12 \p22(Q)^{2} - 16 \p22(Q) b_{1} - 8 b_{2}
	\right) \\ 
	&+\pp112(Q)
	\left(- 4 \p22(Q) - 8 \p22(R)
	\right) \\ 
	&+\pp222(Q)
	\left(- 4 \p11(Q) + 12 \p11(R) + \p12(R) \left(- 12 \p22(R) - 16 b_{1}\right)
	\right) \\ 
	&+\pp112(R)
	\left(8 \p22(Q) + 4 \p22(R)
	\right) \\ 
	&+\pp122(Q)
	\left(8 \p12(Q) + 8 \p12(R) + 12 \p22(R)^{2} + 16 \p22(R) b_{1} + 8 b_{2}
	\right)
	\end{align*}

\bibliographystyle{amsplain}
\bibliography{Bibliography.bib}
\end{document}